# FreeBlend: Advancing Concept Blending with Staged Feedback-Driven Interpolation Diffusion

### v2

@@ Abstract
Concept blending is a promising yet underexplored area in generative models. While recent approaches, such as embedding mixing and latent modification based on structural sketches, have been proposed, they often suffer from incompatible semantic information and discrepancies in shape and appearance. In this work, we introduce FreeBlend, an effective, training-free framework designed to address these challenges. To mitigate cross-modal loss and enhance feature detail, we leverage transferred image embeddings as conditional inputs. The framework employs a stepwise increasing interpolation strategy between latents, progressively adjusting the blending ratio to seamlessly integrate auxiliary features. Additionally, we introduce a feedback-driven mechanism that updates the auxiliary latents in reverse order, facilitating global blending and preventing rigid or unnatural outputs. Extensive experiments demonstrate that our method significantly improves both the semantic coherence and visual quality of blended images, yielding compelling and coherent results.
Machine Learning, ICML
Figure 1: We introduce FreeBlend, a novel, training-free approach that effectively blends concepts to generate new objects through feedback interpolation and auxiliary inference. Unlike previous methods such as TEXTUAL (Melzi et al., 2023), which often suffer from category-specific bias, or Composable Diffusion (Liu et al., 2022), which tends to generate outputs with co-occurring subjects, FreeBlend provides a more robust solution. Additionally, methods like MagicMix (Liew et al., 2022) rely on the structural similarity between the reference and blending concepts, often resulting in unstable blending outcomes. In contrast, FreeBlend consistently produces visually coherent and harmonious blends, setting a new benchmark for state-of-the-art blending techniques.
@@ 1 Introduction
The development of diffusion models (Ho et al., 2020, 2022) has significantly advanced the field of image generation (Rombach et al., 2022; Esser et al., 2023; Brooks et al., 2023; Guo et al., 2023). These models have exhibited impressive abilities of producing novel and appealing images that align with specific instructions, showcasing their potential and flexibility for controlled image synthesis.
@@
Concept blending, as a novel task in the application of diffusion models, involves blending distinct concepts to create a new one that retains the defining characteristics of its components (Fauconnier & Turner, 1998; Ritchie, 2004). It can be viewed as a way of integrating elements from different domains (e.g., cats and dogs) into a novel and meaningful output and creating new objects, scenes, or alterations that are coherent and creatively synthesized. Although current models can generate realistic combinations of individual concepts (Liu et al., 2022; Chefer et al., 2023), more sophisticated concept blending techniques are needed to ensure that the generated images not only blend features but also maintain their logical consistency and visual appeal.
@@
Early methods of concept blending rely on simple vector weighting (Melzi et al., 2023) or hand-crafted feature combinations (Olearo et al., 2024), which involve adding or reassembling text embeddings. These approaches often lead to inaccurate representations and lack correspondence between visual and semantic features due to cross-modal discrepancies. MagicMix (Liew et al., 2022) interpolates the original class latent into another latent space corresponding to a text prompt. While it creatively introduces latent interpolation for mixing, it struggles with shape mismatch issues and lacks flexibility in visual transformations. ConceptLab (Richardson et al., 2024) utilizes VLMs and latent space manipulation, but the constraints imposed during training the additional module limit the flexibility of its application. ATIH (Xiong et al., 2024), based on MagicMix, injects trainable parameters into the denoising process and enforces similarity constraints to harmonize the fusing of texts and images. However, the limitations of its model structure, similar to those of MagicMix, combined with its inability to address mismatched shapes or semantically irrelevant features, hinder its overall blending performance.
In this paper, we introduce FreeBlend, a novel training-free approach to concept blending. At its core, FreeBlend employs a feedback-driven mechanism with latent interpolation to steer the diffusion denoising process. By interpolating auxiliary latents into the initial blending latent and iterating through feedback-driven interactions, the influence of the auxiliary latents gradually diminishes, guiding the blending latent towards the generation of the final image. This integration of blending and denoising processes leverages the strengths of multiple diffusion models to enhance concept blending performance.
@@
Specifically, FreeBlend consists of three core components: transferred unCLIP (Ramesh et al., 2022) image conditions for Stable Diffusion (Rombach et al., 2022), a stepwise increasing interpolation strategy, and a feedback-driven mechanism of denoising process. Instead of using traditional text-based conditions, we employ images generated from text as conditions to guide the generation process via the unCLIP model. This approach provides more precise visual details compared to the vagueness of text-based conditions, reducing the uncertainty caused by cross-modal differences. In addition to this, we divide the denoising process into three stages: the initialization stage, the blending stage, and the refinement stage. At the initialization stage, the pretrained Stable Diffusion model starts with random noise sampled from a unit Normal distribution, which is then denoised under the guidance of unCLIP image condition. At the blending stage, we add noise to the auxiliary latents derived from the condition images to ensure to be in the same period as the blending latent. The blending and auxiliary processes are all denoised simultaneously. In the final refinement stage, only the unCLIP image condition is used to provide additional information, enabling the model to generate images with greater clarity and finer details.
During the blending stage, we set the proportion of the current timestep relative to the total denoising time and use it to adjust the contribution of the auxiliary latents to a blending latent. As denoising progresses, this coefficient decreases, gradually reducing the influence of the auxiliary latents while focusing more on the blending latents, allowing more intricate details to emerge. The feedback-driven mechanism applies this coefficient uniformly, regulating the influence of the auxiliary latents on the blending latents. During the early blending stage, the auxiliary latents enrich the blending latents with information from the original class, and later, they add subtle details without disrupting the overall structure of the image. In addition to qualitative evaluation, we further incorporate quantitative metrics to rigorously assess the generation process, providing a more comprehensive analysis of the model’s performance on the datasets we have generated.
The main contributions of this paper are three-fold:
  • We introduce a novel feedback-driven latent interpolation diffusion model to tackle the concept blending problem. To the best of our knowledge, this is the first training-free method specifically designed for exhaustive concept blending.
  • Our approach incorporates unCLIP to use images as conditions, along with a stepwise increasing interpolation strategy and a feedback-driven denoising process to effectively blend different concepts.
  • Extensive qualitative and quantitative experiments are conducted on several benchmarks, including our proposed blending effect evaluation metric, CLIP-BS, demonstrating that our method achieves state-of-the-art performance in generating blended concepts.
@@ 2 Related Work
Image Editing and Training-free Guidance. Generative models like Stable Diffusion (Rombach et al., 2022) have advanced text-to-image synthesis (Podell et al., 2024; Zhang et al., 2023; Gu et al., 2022) and diffusion models have significantly advanced image editing (Balaji et al., 2022; Couairon et al., 2022; Brooks et al., 2023). DiffEdit (Couairon et al., 2022) and InstructPix2Pix (Brooks et al., 2023) primarily focus on semantic image editing. DiffEdit leverages mask guidance to facilitate intuitive modifications, while InstructPix2Pix enables efficient editing through natural language instructions. Shi et al. (2024) propose DragDiffusion, which improves interactive editing by allowing precise adjustments with handle points and editable regions. In training-free guidance, Feng et al. (2022) introduce structured diffusion guidance to tackle compositional challenges, enhancing the handling of multiple objects. Yu et al. (2023) develop FreeDoM, a training-free method using energy-guided conditional diffusion for dynamic editing with conditions like segmentation maps. Mo et al. (2024) propose FreeControl, offering zero-shot spatial control over pretrained models. These methods enhance the flexibility and precision of diffusion models for effective, retraining-free image editing. Our work builds on this intuition and introduces a novel method for achieving a unique blending style of image editing through training-free guidance.
Concept Composition. The field of diffusion models has seen significant progress in concept composition, where multiple concepts are combined to create complex, multi-faceted images. Liu et al. (2022) introduce a compositional approach that uses multiple models to generate distinct components of an image, effectively addressing challenges related to complex object compositions and their interrelationships. Building on this, Kumari et al. (2023) explore multi-concept customization, enabling the generation of unified, complex images that seamlessly blend multiple concepts. Furthermore, Chefer et al. (2023) propose the Attend-and-Excite method, which improves image accuracy by fine-tuning attention mechanisms, allowing for better semantic alignment between objects in the generated image.
@@
Concept Blending. In contrast, while concept blending holds promise, its practical applications remain more limited and less explored compared to compositional approaches. MagicMix (Liew et al., 2022) addresses semantic mixing while preserving spatial layout, but its use is more restricted due to its shape limitations. Melzi et al. (2023) and Olearo et al. (2024) explore concept blending by manipulating the relationship between different prompts, with Melzi et al. (2023) studying prompt ratios and Olearo et al. (2024) testing various mechanisms for blending text embeddings. Richardson et al. (2024) introduce ConceptLab, using Diffusion Prior models to generate novel concepts within a category through CLIP-based constraints, allowing for the creation of unique hybrid concepts. Xiong et al. (2024) further advance the field with adaptive text-image harmony, merging text and image inputs to generate novel objects while preserving their original layout. While these studies contribute to the domain of concept blending, their overall impact has been more limited compared to concept composition, which has motivated the completion of our work.
3 Methodology
@@ 3.1 Preliminaries
The Stable Diffusion (Rombach et al., 2022) model efficiently operates in the latent space, distinguishing it from previous approaches (Ho et al., 2020; Dhariwal & Nichol, 2021). Specifically, an autoencoder consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$ is trained with a reconstruction objective. Given an image $\mathbf{x}$, the encoder $\mathcal{E}$ maps it to a latent $\mathbf{z}$, and the decoder $\mathcal{D}$ reconstructs the image from the latent, i.e., $\tilde{\mathbf{x}}=\mathcal{D}(\mathbf{z})=\mathcal{D}(\mathcal{E}(\mathbf{x}))$. This way, at each timestep $t$, a noisy latent $\mathbf{z}_{t}$ can be obtained. Beyond the routine training scheme, Stable Diffusion devises conditioning mechanisms to control the synthesized image content with $\mathbf{c}$ representing given additional conditions like text prompts, images, or other measurable properties. The DDPM (Ho et al., 2020) model $\epsilon_{\theta}$ can then be trained via:
@@
where U-Net (Ronneberger et al., 2015) enhanced with self-attention and cross-attention layers is adopted as the denoiser. During training, given a noised latent $\mathbf{z}_{t}$ at timestep $t$ and condition $\mathbf{c}$, denoiser is tasked with predicting the noise $\epsilon_{t}$ added to the current latent.
During inference, a latent $\mathbf{z}_{T}$ is sampled from the standard normal distribution $\mathcal{N}(0,I)$ and the DDPM is used to iteratively remove the noise $\epsilon_{t}=\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ in $\mathbf{z}_{T}$ to produce $\mathbf{z}_{0}$. This process is expressed as:
@@
Figure 2: Overview of our method at the blending stage. Two input images or concepts are encoded into embeddings, mapped to a shared text space via the Linear Prior Converter from unCLIP (Ramesh et al., 2022). These embeddings condition the U-Net: one for downsampling, the other for upsampling. During the blending stage, a blending latent $L_{b}$ initialized with Gaussian noise is processed in the Feedback Interpolation Module, conditioned on image embeddings. Noise $\epsilon$ is added to the embeddings to generate initial auxiliary latents, which are interpolated into $L^{(t)}_{b}$ with an increasing weight $p$. The $L^{(t)}_{a}$ is combined with interpolated latent $L^{\prime(t)}_{b}$ by proportion $p$. All updated $L^{\prime(t)}_{a}$ are refined in the auxiliary inference to retain original features using the text prompt for corresponding categories, and $L^{\prime(t)}_{b}$ is denoised via the blending inference. Finally, the refined $L_{b}$ is passed into the VAE decoder to generate the final blending image.
3.2 unCLIP Image Embeddings for Condition
Previous diffusion models rely on a variety of conditions (Zhan et al., 2024), and traditional concept blending approaches typically use transformed text embeddings (Melzi et al., 2023; Olearo et al., 2024) as inputs to these models. In contrast, our approach utilizes the image embeddings as the sole input, leading to significant improvements in both quality and interpretability. Given that the unCLIP (Ramesh et al., 2022) model, which shares similarities with the DALL-E (Ramesh et al., 2021) mechanism, allows images to serve as conditional inputs, we adopt it as the backbone model for our method.
In order to preprocess $\mathbf{c}$ from the given image modality, we introduce a Linear Prior Converter $\tau_{\theta^{\prime}}$ from unCLIP that maps images from image embeddings into the text embedding space for sampling condition. These embeddings are subsequently used as inputs to the intermediate layers of the U-Net through a cross-attention mechanism, which is defined as
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V,$ (3)
where
$\left\{\begin{array}[]{l}Q=W_{Q}\cdot\varphi(L^{(t)}_{b})\\ K=W_{K}\cdot\tau_{\theta^{\prime}}(\mathbf{c}_{\text{stage}})\\ V=W_{V}\cdot\tau_{\theta^{\prime}}(\mathbf{c}_{\text{stage}}),\\ \end{array}\right.$ (4)
and $\mathbf{c}_{\text{stage}}$ is defined as:
$\mathbf{c}_{\text{stage}}=\begin{cases}\mathbf{c}_{\text{up}}&\text{(for downsampling)}\\ \mathbf{c}_{\text{down}}&\text{(for upsampling)}\end{cases}.$ (5)
Here, $\varphi(L^{(t)}_{b})\in\mathbb{R}^{N\times d_{\epsilon}}$ represents the intermediate (flattened) feature map produced by the U-Net that implements $\epsilon_{\theta}$, and $W_{V}\in\mathbb{R}^{d\times d_{\epsilon}}$, $W_{Q}\in\mathbb{R}^{d\times d_{\tau}}$, and $W_{K}\in\mathbb{R}^{d\times d_{\tau}}$ are learnable projection matrices (Vaswani, 2017; Jaegle et al., 2021). $\mathbf{c}_{\text{up}}$ and $\mathbf{c}_{\text{down}}$ are concept conditions of $\mathbf{c}_{\text{stage}}$, and they are all encoded from original images.
By utilizing the first embedding during the downsampling phase and the second embedding during upsampling, the model is able to seamlessly integrate hybrid features at each training epoch. This methodology mitigates the risk of sequentially introducing two distinct feature sets across epochs, a process that could lead to feature conflicts due to the inherent differences in their nature. Furthermore, a simple averaging of embeddings may result in the attenuation of critical information, such as fine details or spatial positioning. Leveraging the ability to interface with the text embedding space via unCLIP, the model not only captures the semantic essence of the concept but also affords precise control over visual attributes, including shape and color.

### v3

# The First Workshop on Generative and Protective AI for Content Creation

@@ Abstract
Have you ever imagined a creature combining the features of a cat and a car? In this work, we tackle the problem known as concept blending, which represents a fascinating yet underexplored area in generative models. While recent approaches, such as embedding mixing and latent modification based on structural sketches, have been proposed, they often suffer from incompatible semantic information and discrepancies in shape and appearance. In this work, we introduce FreeBlend, an effective, training-free framework designed to address these challenges. To mitigate cross-modal loss and enhance feature details, we leverage transferred image embeddings as conditional inputs. We also design a stepwise increasing interpolation strategy between latents to seamlessly integrate auxiliary features. Moreover, we introduce a feedback-driven mechanism that updates the auxiliary latents in reverse order, facilitating global blending and preventing unnatural outputs. Extensive experiments demonstrate that our method significantly improves both the semantic coherence and visual quality of blended images. The project homepage is available at https://petershen-csworld.github.io/FreeBlend/.
Figure 1: We introduce FreeBlend, a training-free approach that effectively blends two concepts to generate new objects through feedback interpolation and auxiliary inference. This method consistently produces visually coherent and harmonious blends, enabling users to create customized images with diverse combinations of concepts.
@@ 1 Introduction
With the breakneck development of diffusion models, the field of image generation has witnessed rapid progress [rombach2022high, podellsdxl, esser2024scaling, saharia2022photorealistic]. These models exhibit strong instruction-following capabilities and high-fidelity synthesis, achieving remarkable performance across a variety of downstream tasks, including local inpainting [lugmayr2022repaint, corneanu2024latentpaint], personalized image generation [ruiz2023dreambooth, shi2023instantbooth], and more [chen2024anydoor, zhao2023uni, li2024photomaker].
Such advances have paved the way for a spectrum of novel and creative applications. Among them, concept blending stands out as an interesting direction previously overlooked by the community. For example, one could create a creature combining the features of an orange and a teddy bear, resulting in a surprising yet coherent visual concept. More concept pair blends are shown in FreeBlend: Advancing Concept Blending with Staged Feedback-Driven Interpolation Diffusion.
@@
Formally, concept blending involves blending two distinct text concepts to create a new one that retains the defining characteristics of its components [fauconnier1998conceptual, ritchie2004lost]. It can be viewed as a way of integrating elements from different domains into a novel and meaningful output and creating new objects, scenes, or alterations that are coherent and creatively synthesized. Although current models can generate realistic combinations of individual concepts [liu2022compositional, chefer2023attend], more sophisticated concept blending techniques are needed to ensure that the generated images not only blend features but also maintain their semantic consistency and visual appeal.
@@
Early methods of concept blending rely on a simple, “black-box” approach. These methods achieve coarse concept blending by manipulating text embeddings from the input side, which typically involves adding or reassembling the embeddings [melzi2023does, olearo2024blend]. However, these approaches often lead to inaccurate representations and a lack of correspondence between visual and semantic features due to cross-modal discrepancies. MagicMix [liew2022magicmix] interpolates the original class latent into another latent space corresponding to a text prompt. While it creatively introduces latent interpolation for mixing, it struggles with shape-mismatch issues and lacks flexibility in visual transformations. ConceptLab [richardson2024conceptlab] utilizes VLMs and latent space manipulation, but the constraints imposed during training the additional module limit the flexibility of its application. ATIH [xiong2024novel], based on MagicMix, injects trainable parameters into the denoising process and enforces similarity constraints to harmonize the fusing of texts and images. However, the limitations of its model structure, similar to those of MagicMix, combined with its inability to address mismatched shapes or semantically irrelevant features, hinder its overall blending performance.
In this paper, we introduce FreeBlend, a training-free method for concept blending in image generation. As shown in Figure˜2, FreeBlend smoothly interpolates between auxiliary latents by incorporating a feedback-driven mechanism into the diffusion model’s denoising process. As denoising advances, the influence of auxiliary latents diminishes, allowing the blending latent to take greater control, enhancing the blending performance and leveraging the creative power of diffusion models.
@@
Specifically, FreeBlend consists of three core components: transferred unCLIP [ramesh2022hierarchical] image conditions for Stable Diffusion [rombach2022high], a stepwise increasing interpolation strategy, and a feedback-driven mechanism of the denoising process. Instead of using traditional text-based conditions, we employ images generated from text as conditions to guide the generation process via the unCLIP model. This approach reduces the uncertainty caused by cross-modal differences. In addition to this, we divide the denoising process into three stages: the initialization stage, the blending stage, and the refinement stage. At the initialization stage, the pretrained Stable Diffusion model starts with random noise sampled from a standard normal distribution, which is then denoised under the guidance of unCLIP image condition. At the blending stage, we add noise to the auxiliary latents derived from the condition images to ensure to be in the same period as the blending latent. The blending and auxiliary processes are all denoised simultaneously. Ultimately, in the final refinement stage, only the unCLIP image condition is used to provide additional information, enabling the model to generate images with greater clarity and finer details.
The core contributions of this paper are (i) We propose a feedback-driven latent interpolation approach that leverages diffusion models to address the concept blending problem. Our method is designed to be training-free and computationally efficient. (ii) Our approach incorporates unCLIP to use images as conditions, along with a stepwise increasing interpolation strategy and a feedback-driven denoising process to effectively blend different concepts. (iii) We conduct extensive qualitative and quantitative experiments to validate the effectiveness of our proposed method. Multiple evaluation metrics, including our newly introduced blending score metric, CLIP-BS, confirm that our approach achieves state-of-the-art performance in generating blended concepts.
@@ 2 Related Work
Image Editing and Training-free Guidance. Generative models like Stable Diffusion [rombach2022high], Imagen [saharia2022photorealistic], and DALL-E [ramesh2021zero] have advanced text-to-image synthesis [podellsdxl, zhang2023adding, gu2022vector, zhou2023shifted] and diffusion models have significantly advanced image editing [couairon2023zero, couairon2023zero, balaji2022ediff, couairon2022diffedit, brooks2023instructpix2pix, shi2024dragdiffusion], image interpolation [wang2023interpolating, qiyuan2024aid], style transfer [hamazaspyan2023diffusion, wang2023stylediffusion], and 3D generation [xiang2024structured]. DiffEdit [couairon2022diffedit] and InstructPix2Pix [brooks2023instructpix2pix] primarily focus on semantic image editing. DiffEdit leverages mask guidance to facilitate intuitive modifications, while InstructPix2Pix enables efficient editing through natural language instructions. In training-free guidance, Structure Diffusion [feng2022training] introduces structured diffusion guidance to tackle compositional challenges, enhancing the handling of multiple objects. FreeDoM [yu2023freedom] is a training-free method using energy-guided conditional diffusion for dynamic editing with conditions like segmentation maps. FreeControl [mo2024freecontrol] offers zero-shot spatial control over pretrained models. These methods enhance the flexibility and precision of diffusion models for effective, retraining-free image editing. Our work builds on this intuition and introduces a novel method for achieving a unique blending style of image editing through training-free guidance.
@@
Concept Composition and Blending. The field of concept composition has made significant advancements, with the combination of multiple concepts to generate complex, multi-faceted images [liu2021learning, liu2022compositional, wang2024compositional, kumari2023multi, chefer2023attend]. Composable Diffusion [liu2022compositional] introduces a compositional approach that utilizes multiple models to generate distinct components of an image, effectively addressing challenges associated with complex object compositions and their interrelationships. Custom Diffusion [kumari2023multi] explores multi-concept customization, enabling the generation of unified, intricate images that seamlessly blend multiple concepts. In contrast, while concept blending holds substantial promise, its practical applications remain more limited and underexplored when compared to compositional approaches. MagicMix [liew2022magicmix] addresses semantic mixing while preserving spatial layout, though its use is more restricted due to shape limitations. Melzi et al. [melzi2023does] and Olearo et al. [olearo2024blend] investigate concept blending by manipulating the relationships between different prompts, with the former focusing on prompt ratios and the latter experimenting with various mechanisms for blending text embeddings. ConceptLab [richardson2024conceptlab] leverages Diffusion Prior models to generate novel concepts within a category through CLIP-based constraints, enabling the creation of unique hybrid concepts. ATIH [xiong2024novel] further advances the field by introducing adaptive text-image harmony, merging text and image inputs to generate novel objects while maintaining their original layout. Although these studies contribute to the domain of concept blending, their overall impact has been more limited compared to concept composition, which has inspired the completion of our work.
3 Method
@@ 3.1 Preliminaries
Our method is based on Stable Diffusion, a prominent application of the latent diffusion model [rombach2022high]. It operates within a compressed latent space, improving sampling and denoising efficiency compared to DDPM [ho2020denoising] and DDIM [dhariwal2021diffusion]. A VAE [kingma2013auto] is used to compress image representations, enhancing computational efficiency. During training, at timestep $t$, a noisy latent vector $\mathbf{z}_{t}$ is generated by adding noise to the latent representation $\mathbf{z}_{0}$ (the original clean image latent). This noisy latent is progressively denoised over several timesteps to recover the original image.
Besides, Stable Diffusion employs classifier-free guidance  [ho2021classifier] to control the synthesized image content with $\mathbf{c}$ representing given additional conditions like text prompts, images, or other customized properties. The model $\epsilon_{\theta}$ could be trained via the following objective:
@@
where $\epsilon_{\theta}$, typically implemented by U-Net [ronneberger2015u] or DiT [peebles2023scalable], represents the denoiser.
During inference, a latent $\mathbf{z}_{t}$ is sampled from the standard normal distribution $\mathcal{N}(0,I)$ and the trained denoiser is used to iteratively remove the noise $\epsilon_{t}=\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ in $\mathbf{z}_{t}$ to produce $\mathbf{z}_{0}$. This process is expressed as:
@@
Figure 2: Overview of our method. Two input images, generated by Stable Diffusion with respective concepts, are encoded into CLIP embeddings and mapped to a shared text space via the Linear Prior Converter from unCLIP [ramesh2022hierarchical]. These embeddings condition the U-Net, one for downsampling and the other for upsampling. The blending latent $L_{b}$ is initialized with Gaussian noise and processed during initialization. The module within the dashed box is used only in the blending stage. Noise $\epsilon$ is added to the image embeddings to generate initial auxiliary latents, which are interpolated into $L^{(t)}_{b}$ for feedback. The latent $L^{(t)}_{a}$ is combined with $L^{\prime(t)}_{b}$ by proportion pp. Updated latents $L^{\prime(t)}_{a}$ are refined in auxiliary inference using unCLIP embeddings to preserve original features, while $L^{\prime(t)}_{b}$ is denoised in the blending inference. Finally, the blending latent is refined and passed to the VAE decoder to generate the final image.
3.2 Concept Blending with unCLIP Condition
In multimedia production, the innovation of novel entities is crucial, particularly in the development of anthropomorphic characters. A prevalent method for conceptual innovation involves the fusion of distinct ideas. For instance, numerous animated characters are conceived by combining elements from various sources, such as the "Iron Man" suit and a traditional Japanese samurai’s armor.
In this section, we present a novel image generation task called concept blending. This task aims to merge two distinct text concepts and generate blended images that retain the original shape, color, texture, and other relevant details. To achieve this task, we first attempt to directly use texts as condition [melzi2023does, olearo2024blend], but due to the low quality and visual uncertainty due to the gap of vision and text, we decide to choose images as condition [zhan2024conditional, ye2023ip, mou2024t2i].
We adopt the Linear Prior Converter from unCLIP [ramesh2022hierarchical], which, like DALL-E [ramesh2021zero], enables images as conditional inputs. This converter maps image embeddings to the text embedding space, and the resulting embeddings are fed into the U-Net’s intermediate layers via cross-attention, guiding the denoising process across all stages. While mapping image embeddings to the text space, the transformed embeddings preserve visual-like information, similar to textual inversion [gal2022image], providing more deterministic features and richer details than the original text embeddings.
Our model uses two types of embeddings at different phases: one during the downsampling phase and another during the upsampling phase. Simply averaging the embeddings [melzi2023does] could lead to the loss of essential details, such as fine textures or spatial positioning. By interacting with the text embedding space through unCLIP, the model captures the semantic essence of the concepts while maintaining precise control over visual attributes like shape and color.